A block of mass m = 7.40 kg is released from rest
Answer #1

KE1 + PE1 = KE2 + PE2

0 + 7.40*9.8*5.20 = 1/2*7.40*V^2 + 7.40*9.8*3.20

=> VB = 6.26 m/s

0 + 7.40*9.8*5.20 = 1/2*7.40*V^2 + 7.40*9.8*2

=> VC = 7.91 m/s

b) work = F*X = 7.40*9.8*3 = 232 J